# Digital Twin for Wind Energy: Latest updates from the NorthWind project Аннотация

### v1

@@ 2 Digital twin
Following an extensive literature survey [2] ‘‘A digital twin is defined as a virtual representation of a physical asset or process enabled through data and simulators for real-time prediction, optimization, monitoring, control and improved decision-making. "
To distinguish between the capabilities of digital twins, they can be classified on a scale from 0-5 (0-standalone, 1-descriptive, 2-diagnostic, 3-predictive, 4-prescriptive, 5-autonomous) [1]. The standalone digital twin is a virtual representation disconnected from the real asset and can be used during the design and planning stage. The descriptive digital twin describes the current state of the asset based on a real-time data stream. At the diagnostic level, the digital twin can be used to identify faults based on the current and historic asset state. In contrast to previous levels, the predictive digital twin also possesses the capability to forecast future asset states. The prescriptive digital twin provides recommendations to the user based on uncertainty quantification and risk analysis. Finally, the autonomous digital twin closes the loop as it optimizes the asset through bidirectional data exchange where the asset updates the digital twin and the digital twin controls the asset. More detailed explanations of the capability levels in the context of wind energy can be found in [1]. The article concludes that the primary challenges impeding the development and widespread adoption of the technology fall into three broad categories: data-related, model-related, and industrial acceptance-related challenges. The subsequent focus within the project has been on addressing these challenges, which are elaborated upon in the following section.
@@ 3.1.2 Synthetic data generation
The data for the Bessaker wind farm were generated using a HARMONIE-SIMRA multiscale wind modeling system ([6],[7]). horizontal resolution wind forecast data from the HARMONIE model was used to force the SIMRA model which, in turn, generated a wind field at a fine horizontal resolution of over a domain of . The simulation sample data can be seen in Figure 3. This data were used to train a Generative Adversarial Network model as explained in Section 3.3.3.
@@ 3.2.2 Data integration framework
Commencing from the descriptive level (level 1), digital twins are envisioned to synchronize seamlessly with their physical counterparts, providing an intricate representation of their current and historical states. This entails, firstly, aggregating data on the asset and its surroundings from various sources. Secondly, given the often sparse spatial and temporal dimensions of data, alongside the inability to measure certain parameters directly, models are indispensable for interpolating, extrapolating, and inferring data, including future values. Thirdly, digital twins necessitate the ability to present both data and analysis outcomes in an accessible manner across diverse clients and platforms.
@@
This data integration framework was demonstrated successful implementation in constructing digital twins for an onshore wind farm [4] and an offshore wind turbine [3].
@@ 3.3.1 Component based reduced order models
The text describes a model of an offshore wind turbine (OWT) jacket foundation and tower structure. The structure consists of tubular steel beams and is attached to the seabed at z=-27.5m. The jacket is m high, and the tower is 86.4m high (the waterline is at z=0). Offshore jacket structures are often modeled using beam finite elements only. However, more refined modeling in certain regions, such as tubular joints, may be performed in order to improve numerical accuracy. Furthermore, for fatigue or ultimate limit-state post processing according to industry recommended practices, the conservatism of the computations may be reduced when shell models are considered. In the example here, the tubular joints between chords and braces are modeled using shell finite elements. Furthermore, eight of the tubular beam-element joints are replaced by three-dimensional shell-element reduced-basis components.
Each component has six ports, each with three translational and three rotational degrees of freedom. The degrees of freedom on the circular boundaries of the shell elements are condensed to a single point using explicit constraint equations. The finite-element discretization of each component consists of 26241 first-order triangular (three-noded) and quadrilateral (four-noded) linear shell finite elements. The resulting number of degrees of freedom is . Each component is parameterized by the wall thickness of the top and bottom braces. The parameter enters into the equations as the thickness parameter in the shell finite element stiffness matrices. However, note that, due to the automatic affinization procedure, no detailed knowledge of the parameterization is needed. For the discretization of the remaining jacket structure, 102 first-order (two-noded) beam elements were utilized. A single beam element was employed between each joint, while the tower comprises multiple beam elements of varying diameters to approximate a conical shape. This section of the structure does not incorporate any parameters.
In the offline stage, Aroma was utilized to generate the component using 50 snapshots for each of the six ports and each of the six degrees of freedom on each port. The sampling parameters are distributed according to the roots of the Legendre polynomials. Aroma invokes Sesam Sestra as the full-order solver.
@@
Utilizing OpenFOAM’s SIMPLE algorithm, an ensemble of snapshots was gathered, starting with the upstream airfoil in free-stream. Solutions were obtained for three Reynolds numbers: , , and . These snapshots served as the basis for formulating reduced bases for velocity and pressure, employing Galerkin projection. The free-stream ROM approximation was then utilized to establish accurate boundary conditions, replicating the downstream simulation. Subsequently, for the now-perturbed flow, an additional set of snapshots was collected, expanding the original sampling space. A new ROM was constructed, capable of approximating both upstream and downstream elements.
@@ 3.3.2 Windfarm layout optimization
In this work, the utilization of the open-source Julia code (FLOWFarm.jl, [14]) is demonstrated. The code is applied to the Unitech Zefyros virtual wind farm positioned at project area 3 for the Utsira Nord tender announced by the Norwegian Government in 2023. The requirement in the tender specified a minimum of of installed capacity energy density for the farm in this area, which spans . This leaves approximately of surplus area in relation to the government’s energy density requirements. Assuming that the potential expansion into this surplus area is contingent on improvements in the network connection capacity, the entire area could be utilized at the specified density, resulting in an installed capacity of . Opting for the NREL turbine as a reference, this translates to 41 turbines (or 125 turbines for turbines). While acknowledging that not the entire area would realistically be utilized, the full area is employed for the purpose of comparing the two wake model codes. The process begins with the extraction of a wind rose for 100-m winds from ERA5 data, collected for every sixth hour from the year 2000 through December 2022. These data are sourced from a point located at 4.5E and 59.1N just offshore the west coast of Norway. To align with the chosen NREL turbine, the wind speed level is extended from to using a power law approach with a shear exponent of 0.15. The windrose bins are divided into 10 degrees (36 directional bins), with the average wind speed for each directional bin applied during optimization.
@@ 3.3.3 GANs
GANs consist of two antagonistic neural networks—the generator and the discriminator. The generator, given an input, produces an output, while the discriminator assesses whether the output appears genuine or artificial. Through joint training, the generator becomes proficient in generating results that closely resemble real-world scenarios. In the current study, GANs are trained by inputting coarse-scale simulations corresponding to a horizontal resolution of and compelling the network to generate a high-resolution field with a resolution of . The resolution is derived by applying a nearest neighbor algorithm to the synthesized data with a resolution. What distinguishes this work from prior efforts in image resolution upscaling is the incorporation of a cost function that adheres to the principles of physics. More details of the methodology can be found in [15].
@@ 4.1 Reverse engineered data
Designing a wind turbine control system is a multifaceted task considering numerous variables to ensure the turbine is efficient, safe, and reliable. A simplified control system has been developed primarily centered around wind speed, rotor speed, and blade pitch. The development of this simplified control system involved a series of Aerodyn simulations, with pre-described airfoil properties. The starting point for this control system was the SWT 2.3-93 wind turbine. Throughout the Aerodyn simulations, adjustments were made to both the turbine speed and blade pitch to achieve a power curve of the SWT 2.3-83 wind turbine. Utilizing the developed control system and blade aerodynamic geometry, Aerodyn simulations of the turbine were conducted. The simulated power curve compared with the actual power curve is presented and shown in Figure. The reconstructed turbine geometry showed a good comparison with the anticipated power output across varying wind speeds. Through an iterative process, turbine geometry was successfully built that closely reflected the expected performance characteristics. The studies showed that the reverse engineering approach can effectively be used to design wind turbine geometry, which signifies a significant milestone in developing a digital twin for the turbine. However, while these results are promising, further validation and refinement are essential for enhancing the accuracy and applicability of this methodology, particularly under diverse operational conditions and environmental factors.
@@ 4.5 GANs enabled microscale windflow simulations
Figure 14 displays the outcomes of the GANs-based reconstruction of the high-resolution wind field. The figure comprises four subfigures (LR, HR, RL, SR). LR represents a coarse-scale simulation (800m800m) of the wind field in complex terrain, while HR corresponds to a horizontal resolution of 200m200m. TL denotes the trilinear interpolation of the LR field on a mesh corresponding to HR, serving as the ground truth. SR illustrates the upscaled resolution of the wind field achieved through the GANs network. It is clear that the SR field bears a striking resemblance to the one obtained through trilinear interpolation. The integration of GANs facilitates a computational speedup of up to 100 times without compromising accuracy. Additional details about this work are available in our archived article [15].
@@ 5 Conclusion and Future Work
In the current work, the authors present some of the highlights of the research conducted within the NorthWind project. Due to page limitations, not all required details could be provided. However, citations to papers where more details can be found are included. The presented work can be summarized as follows
@@
Generative Adversarial Networks (GANs) were presented for upscaling the resolution of wind fields, attaining the accuracy of a 200m200m simulation at the computational cost of 800m800m resolution simulations, resulting in a 100x speed-up of high-resolution wind simulations in complex terrain.
@@
Future work envisions enhanced utilization of the asset information model in numerical simulations and control processes, optimizing its effectiveness. To achieve this, the adoption of evolutionary algorithms for reverse engineering is proposed, aiming for a more efficient alternative to the current brute-force approach. The windfarm layout optimization process is anticipated to evolve by incorporating additional factors and constraints, further refining the models for enhanced decision-making. Additionally, the development of Component-ROMS for fluids and fluid-structure interactions is identified as a key focus for future research. It is important to highlight that although the authors presented high fidelity wake simulation data, it has not yet been used for building Component-ROMS. Once such a component-ROM is operational, it can be utilized within the wind farm layout optimization work presented in this study. In the future, efforts will be directed towards expanding the capabilities of GANS to model the marine boundary layer, with a particular focus on understanding the impacts of stratification. Additionally, there is a plan to seamlessly integrate the data flow framework, asset information model, simulations, and control algorithms. This comprehensive integration is expected to yield an intuitive human-machine interface, offering a robust platform for efficient decision-making. This integration process will improve the overall functionality and user experience, thus contributing to advancements in wind energy technology.
@@ Acknowledgments
This publication has been prepared as part of NorthWind (Norwegian Research Centre on Wind Energy) co-financed by the Research Council of Norway (project code 321954), industry, and research partners. Read more at www.northwindresearch.no. Data for the project realization was provided by SINTEF Energy Research.
@@
- [1] Stadtmann, Florian, Rasheed, Adil, Kvamsdal, Trond, Johannessen, Kjetil André, San, Omer, Kölle, Konstanze, Tande, John Olav, Barstad, Idar, Benhamou, Alexis, Brathaug, Thomas, Christiansen, Tore, Firle, Anouk-Letizia, Fjeldly, Alexander, Frøyd, Lars, Gleim, Alexander, Høiberget, Alexander, Meissner, Catherine, Nygård, Guttorm, Olsen, Jørgen, Paulshus, Håvard, Rasmussen, Tore, Rishoff, Elling, Scibilia, Francesco and Skogås, John Olav. ‘‘Digital Twins in Wind Energy: Emerging Technologies and Industry-Informed Future Directions.’’ IEEE Access Vol. 11 (2023): pp. 110762–110795.
@@
- [12] Eivind, Fonn, van Brummelen, Harald, Kvamsdal, T and Rasheed, Adil. ‘‘Fast divergence-conforming reduced basis methods for steady Navier–Stokes flow.’’ Computer Methods in Applied Mechanics and Engineering, Volume 346, (2019), pp. 486-512, ISSN 0045-7825.
- [13] Quarteroni, Alfio and Rozza, Gianluigi (eds.). Reduced Order Methods for Modeling and Computational Reduction. Springer International Publishing, Cham (2014).

### v2

@@ 2 Digital twin
Following an extensive literature survey [2] ‘‘A digital twin is defined as a virtual representation of a physical asset or process enabled through data and simulators for real-time prediction, optimization, monitoring, control and improved decision-making.’’ The concept is depicted in Figure 1. The physical asset, in this case, a wind farm, is shown on the top right side. The state of the asset is measured in real-time through various sensors, resulting in a large data stream from the asset. Still, the data are coarse in spatial-temporal resolution and do not describe the physical behavior and future states of the asset. By exploiting physical models, the data can be complemented to instill physical realism into a virtual representation of the physical asset: the digital twin. This virtual representation can then be used to perform decision-making or even control the asset autonomously. Real-time data exchange in Figure 1 is marked with green arrows, while gray arrows and boxes describe the offline operations. Long-term decisions and recommendations can be provided by exploring hypothetical scenarios, e.g. for risk assessment or uncertainty quantification. This non-real-time application of the digital twin framework is also referred to as digital sibling. Note that while in the figure the virtual representation is labeled digital twin, data processing and analysis, physical models, decision support, and control are all implemented into the digital twin framework, and are therefore part of the digital twin.
To distinguish between the capabilities of digital twins (Figure 2), they can be classified on a scale from 0-5 (0-standalone, 1-descriptive, 2-diagnostic, 3-predictive, 4-prescriptive, 5-autonomous) [1] . The standalone digital twin is a virtual representation disconnected from the real asset and can be used during the design and planning stage. The descriptive digital twin describes the current state of the asset based on a real-time data stream. At the diagnostic level, the digital twin can be used to identify faults based on the current and historic asset state. In contrast to previous levels, the predictive digital twin also possesses the capability to forecast future asset states. The prescriptive digital twin provides recommendations to the user based on uncertainty quantification and risk analysis. Finally, the autonomous digital twin closes the loop as it optimizes the asset through bidirectional data exchange where the asset updates the digital twin and the digital twin controls the asset. More detailed explanations of the capability levels in the context of wind energy can be found in [1]. The paper concludes that the primary challenges impeding the development and widespread adoption of the technology fall into three broad categories: data-, model-, and industrial acceptance-related challenges. The subsequent focus within the project has been on addressing these challenges, see the following section.
@@ 3.1.2 Synthetic data generation
The data for the Bessaker wind farm were generated using a HARMONIE-SIMRA multiscale wind modeling system ([6],[7]). Wind forecast data from the HARMONIE model with a horizontal resolution was used to force the SIMRA model which, in turn, generated a wind field at a fine horizontal resolution of over a domain of . The simulation sample data can be seen in Figure 3. This data were used to train a Generative Adversarial Network model as explained in Section 3.3.3.
@@ 3.2.2 Data integration framework
Commencing from the descriptive level (capability level 1), digital twins are envisioned to synchronize seamlessly with their physical counterparts, providing an intricate representation of their current and historical states. This entails, firstly, aggregating data on the asset and its surroundings from various sources. Secondly, given the often sparse spatial and temporal dimensions of data, alongside the inability to measure certain parameters directly, models are indispensable for interpolating, extrapolating, and inferring data, including future values. Thirdly, digital twins necessitate the ability to present both data and analysis outcomes in an accessible manner across diverse clients and platforms.
@@ 3.3.1 Component based reduced order models
The text describes a model of an offshore wind turbine (OWT) jacket foundation and tower structure. The structure consists of tubular steel beams and is attached to the seabed at . The jacket is high, and the tower is high (the waterline is at ). Offshore jacket structures are often modeled using beam finite elements only. However, more refined modeling in certain regions, such as tubular joints, may be performed in order to improve numerical accuracy. Furthermore, for fatigue or ultimate limit-state post processing according to industry recommended practices, the conservatism of the computations may be reduced when shell models are considered. In the example here, the tubular joints between chords and braces are modeled using shell finite elements. Furthermore, eight of the tubular beam-element joints are replaced by three-dimensional shell-element reduced-basis components.
Each component has six ports, each with three translational and three rotational degrees of freedom. The degrees of freedom on the circular boundaries of the shell elements are condensed to a single point using explicit constraint equations. The finite-element discretization of each component consists of 26241 first-order triangular (three-noded) and quadrilateral (four-noded) linear shell finite elements. The resulting number of degrees of freedom is . Each component is parameterized by the wall thickness of the top and bottom braces. The parameter enters into the equations as the thickness parameter in the shell finite element stiffness matrices. However, note that, due to the automatic affinization procedure, no detailed knowledge of the parameterization is needed. For the discretization of the remaining jacket structure, 102 third-order (two nodes with six degrees of freedom at each node) beam elements SESAM Sestra were utilized. A single beam element was employed between each joint, while the tower comprises multiple beam elements of varying diameters to approximate a conical shape. This section of the structure does not incorporate any parameters.
In the offline stage, Aroma was utilized to generate the component using 50 snapshots for each of the six ports and each of the six degrees of freedom on each port. The sampling parameters are distributed according to the roots of the Legendre polynomials. Aroma invoked the Sesam Sestra as the full-order solver.
@@
Utilizing OpenFOAM’s SIMPLE algorithm, an ensemble of snapshots was gathered, starting with the upstream airfoil in free-stream. Solutions were obtained for three Reynolds numbers:. These snapshots served as the basis for formulating reduced bases for velocity and pressure, employing Galerkin projection. The free-stream ROM approximation was then utilized to establish accurate boundary conditions, replicating the downstream simulation. Subsequently, for the now-perturbed flow, an additional set of snapshots was collected, expanding the original sampling space. A new ROM was constructed, capable of approximating both upstream and downstream elements.
@@ 3.3.2 Windfarm layout optimization
In this work, the utilization of the open-source Julia code (FLOWFarm.jl, [14]) is demonstrated. The code is applied to the Unitech Zefyros virtual wind farm positioned at project area 3 for the Utsira Nord tender announced by the Norwegian Government in 2023. The requirement in the tender specified a minimum of of installed capacity energy density for the farm in this area, which spans . This leaves approximately of surplus area in relation to the government’s energy density requirements. Assuming that the potential expansion into this surplus area is contingent on improvements in the network connection capacity, the entire area could be utilized at the specified density, resulting in an installed capacity of . Opting for the NREL turbine as a reference, this translates to 41 turbines (or 125 turbines for turbines). While acknowledging that not the entire area would realistically be utilized, the full area is employed for the purpose of comparing the two wake model codes. The process begins with the extraction of a wind rose for 100-m winds from ERA5 data, collected for every sixth hour from the year 2000 through December 2022. These data are sourced from a point located at and just offshore the west coast of Norway. To align with the chosen NREL turbine, the wind speed level is extended from to using a power law approach with a shear exponent of 0.15. The windrose bins are divided into (36 directional bins), with the average wind speed for each directional bin applied during optimization.
@@ 3.3.3 GANs
GANs consist of two antagonistic neural networks—the generator (G) and the discriminator (D) (see Figure 7). The generator, given an input, produces an output, while the discriminator assesses whether the output appears genuine or artificial. Through joint training, the generator becomes proficient in generating results that closely resemble real-world scenarios. In the current study, GANs are trained by inputting coarse-scale simulations corresponding to a horizontal resolution of and compelling the network to generate a high-resolution field with a resolution of . The resolution is derived by applying a nearest neighbor algorithm to the synthesized data with a resolution. What distinguishes this work from prior efforts in image resolution upscaling is the incorporation of a cost function that adheres to the principles of physics. More details of the methodology can be found in [15].
@@ 4.1 Reverse engineered data
Designing a wind turbine control system is a multifaceted task considering numerous variables to ensure the turbine is efficient, safe, and reliable. A simplified control system has been developed primarily centered around wind speed, rotor speed, and blade pitch. The development of this simplified control system involved a series of Aerodyn simulations, with pre-described airfoil properties. The starting point for this control system was the SWT 2.3-93 wind turbine. Throughout the Aerodyn simulations, adjustments were made to both the turbine speed and blade pitch to achieve a power curve of the SWT 2.3-83 wind turbine. Utilizing the developed control system and blade aerodynamic geometry, Aerodyn simulations of the turbine were conducted. The simulated power curve compared with the actual power curve is presented and shown in Figure 8. The reconstructed turbine geometry showed a good comparison with the anticipated power output across varying wind speeds. Through an iterative process, turbine geometry was successfully built that closely reflected the expected performance characteristics. The studies showed that the reverse engineering approach can effectively be used to design wind turbine geometry, which signifies a significant milestone in developing a digital twin for the turbine. However, while these results are promising, further validation and refinement are essential for enhancing the accuracy and applicability of this methodology, particularly under diverse operational conditions and environmental factors.
@@ 4.5 GANs enabled microscale windflow simulations
Figure 14 displays the outcomes of the GANs-based reconstruction of the high-resolution wind field. The figure comprises four subfigures (LR, HR, RL, SR). LR represents a coarse-scale simulation of the wind field in complex terrain, while HR corresponds to a horizontal resolution of . TL denotes the trilinear interpolation of the LR field on a mesh corresponding to HR, serving as the ground truth. SR illustrates the upscaled resolution of the wind field achieved through the GANs network. It is clear that the SR field bears a striking resemblance to the one obtained through trilinear interpolation. The integration of GANs facilitates a computational speedup of up to 100 times without compromising accuracy. Additional details about this work are available in our archived article [15].
@@ 5 Conclusion and Future Work
In the current work, the authors present some of the highlights of the research conducted within the NorthWind project. The presented work can be summarized as follows
@@
Generative Adversarial Networks (GANs) were presented for upscaling the resolution of wind fields, attaining the accuracy of a simulation at the computational cost of resolution simulations, resulting in a 100x speed-up of high-resolution wind simulations in complex terrain.
@@
Future work envisions enhanced utilization of the asset information model in numerical simulations and control processes, optimizing its effectiveness. To achieve this, the adoption of evolutionary algorithms for reverse engineering is proposed, aiming for a more efficient alternative to the current brute-force approach. The windfarm layout optimization process is anticipated to evolve by incorporating additional factors and constraints, further refining the models for enhanced decision-making. Additionally, the development of Component-ROMS for fluids and fluid-structure interactions is identified as a key focus for future research. It is important to highlight that although the authors presented high fidelity wake simulation data, it has not yet been used for building Component-ROMS. Once such a component-ROM is operational, it can be utilized within the wind farm layout optimization work presented in this study.
In the future, efforts will be directed towards expanding the capabilities of GANS to model the marine boundary layer, with a particular focus on understanding the impacts of stratification. Additionally, there is a plan to seamlessly integrate the data flow framework, asset information model, simulations, and control algorithms. This comprehensive integration is expected to yield an intuitive human-machine interface, offering a robust platform for efficient decision-making. This integration process will improve the overall functionality and user experience, thus contributing to advancements in wind energy technology.
@@ Acknowledgments
This publication has been prepared as part of NorthWind (Norwegian Research Centre on Wind Energy) co-financed by the Research Council of Norway (project code 321954), industry, and research partners. More details are given at www.northwindresearch.no. Data for the project realization was provided by SINTEF Energy Research.
@@
- [1] Stadtmann, Florian, Rasheed, Adil, Kvamsdal, Trond, Johannessen, Kjetil André, et al. ‘‘Digital Twins in Wind Energy: Emerging Technologies and Industry-Informed Future Directions.’’ IEEE Access Vol. 11 (2023): pp. 110762–110795.
@@
- [12] Eivind, Fonn, van Brummelen, Harald, Kvamsdal, Trond and Rasheed, Adil. ‘‘Fast divergence-conforming reduced basis methods for steady Navier–Stokes flow.’’ Computer Methods in Applied Mechanics and Engineering, Volume 346, (2019), pp. 486-512, ISSN 0045-7825.
- [13] Quarteroni, Alfio and Rozza, Gianluigi. ‘‘Reduced Order Methods for Modeling and Computational Reduction’’. Springer International Publishing, Cham (2014).